# PointNetPGAP-SLC: A 3D LiDAR-based Place Recognition Approach with Segment-level Consistency Training for Mobile Robots in Horticulture

### v1

@@ Abstract
This paper addresses robotic place recognition in horticultural environments using 3D-LiDAR technology and deep learning. Three main contributions are proposed: (i) a novel model called PointNetPGAP, which combines a global average pooling aggregator and a pairwise feature interaction aggregator; (ii) a Segment-Level Consistency (SLC) model, used only during training, with the goal of augmenting the contrastive loss with a context-specific training signal to enhance descriptors; and (iii) a novel dataset named HORTO-3DLM featuring sequences from orchards and strawberry plantations. The experimental evaluation, conducted on the new HORTO-3DLM dataset, compares PointNetPGAP at the sequence- and segment-level with state-of-the-art (SOTA) models, including OverlapTransformer, PointNetVLAD, and LOGG3D. Additionally, all models were trained and evaluated using the SLC. Empirical results obtained through a cross-validation evaluation protocol demonstrate the superiority of PointNetPGAP compared to existing SOTA models. PointNetPGAP emerges as the best model in retrieving the top-1 candidate, outperforming PointNetVLAD (the second-best model). Moreover, when comparing the impact of training with the SLC model, performance increased on four out of the five evaluated models, indicating that adding a context-specific signal to the contrastive loss leads to improved descriptors. The dataset and the code will be made publicly available at https://github.com/Cybonic/PointNetPGAP-SLC.git
@@ I INTRODUCTION
Autonomous robots have become more common in modern agriculture, where they have been used for various tasks such as the application of pesticides and fertilizers, infrastructure inspection (e.g., fences and irrigation systems), and fruit and vegetable harvesting [1]. Agricultural robots require precise and accurate localization systems to operate reliably. In horticultural environments, plantations may be located within polytunnels, greenhouses or in places with lack of GNSS coverage. Therefore, in such environments, a GNSS-based localization systems may not be a reliable localization source. Alternatively, most localization systems rely on map-based approaches, where Simultaneous Localization and Mapping (SLAM) can be used to generate or update maps [2].
As part of SLAM systems, or as a standalone global localization approach, 3D LiDAR-based place recognition has seen significant advancements, driven primarily by the autonomous driving community, with the goal of achieving long-term localization [3, 4]. Despite recent efforts to adapt these methods to natural [5] and agricultural environments [6], major challenges still remain largely unresolved.
This work addresses the 3D LiDAR place recognition problem for robots operating in horticultural environments. Horticultural crops present unique challenges due to varying planting arrangements, both between different plant species and within the same species. Factors such as inter-row spacing, plant arrangement and density affect the rows’ permeability to LiDAR scans, allowing beams to penetrate multiple rows. This phenomenon generates sparse and highly overlapping scans with poor geometries, which leads to inter-row descriptor ambiguity.
To cope with these challenges, we propose three key contributions:
(i) PointNetPGAP: A lightweight 3D LiDAR-based place recognition model that utilizes PointNet [7] for local feature extraction and aggregates the local features by combining a global average pooling with another aggregator that explores pairwise feature interactions into a single descriptor.
(ii) Segment-Level Consistency (SLC) model: A model introduced during training to address the inter-row descriptor ambiguity problem. Specifically, we split the crops in various segments, and use the SLC model to provide a training signal, in addition to the contrastive loss, that differentiates scans from distinct rows/segments, even if they are highly overlapped. The SLC model consists of a Multi-layer Perceptron (MLP) which predicts the segment class for each descriptor and computes the loss based on the corresponding ground-truth. By relying on the descriptors as inputs, the SLC model is easily integrated with any place recognition model that outputs descriptors.
(iii) Horticulture 3D Localization and Mapping (HORTO-3DLM) Dataset: A new dataset from horticultural environments comprising data from three orchards and one strawberry plantation using a 3D LiDAR sensor and RTK-GPS mounted on a mobile platform, as shown in Fig. 1. Although, in this work, this dataset is primarily employed for place recognition, it can also be utilized for other localization and mapping tasks.
We benchmark PointNetPGAP against state-of-the-art (SOTA) models, namely OverlapTransformer [8], PointNetVLAD [9], and LOGG3D-Net [10] on the HORTO-3DLM dataset using a cross-validation evaluation protocol. The results demonstrate that PointNetPGAP offers a more effective and efficient solution compared to the aforementioned SOTA models. Furthermore, we train all models with the SLC approach and report the results both at the sequence- and segment-level. Notably, all models except OverlapTransformer exhibit performance improvements, which is primarily driven by an increase in performance in segments with higher ambiguity.
The remainder of this paper is organized as follows. In Section II, we discuss related work. Section III outlines the proposed method. The experimental evaluation and the results are presented and discussed in Section IV, while the concluding remarks are provided in Section V.
@@
Figure 2: Top: Retrieval framework using PointNetPGAP to generate global descriptors. The framework comprises the following key stages: T1) An input scan $P$ is processed by PointNet to extract local features $F$; T2) These local features are aggregated into a global descriptor $D$ by concatenating the global average representation of local features with pairwise feature interactions into a single representation; T3) The global descriptor queries the database to retrieve the top-$k$ most similar places. Bottom: Training scheme with the following stages: B1) A training tuple consisting of scans, including an anchor-positive pair and $m$ negatives, is input into the model, which generates the corresponding descriptors; B2) In the SLC model, the descriptors are input into an MLP that predicts the corresponding segment classes. The loss is computed using the Negative Log-likelihood loss based on the predictions and ground-truth segment labels; B3) The LazyTriplet loss is calculated based on the distances between the anchor-positive pair and the anchor-negatives; B4) Finally, both losses are combined using a weight $\alpha$.
@@ II RELATED WORK
Place recognition has been the subject of much research over the past decade, particularly in recent years, with a focus on 3D-LiDAR approaches [3]. LiDAR sensory data has been recognized as a relevant modality to achieve robustness in place recognition, primarily due to its relative invariance to challenging variations such as different seasonal or lighting conditions, which are important in horticultural robotics.
@@
Regarding aggregation strategies, most 3D LiDAR place recognition frameworks rely on first-order aggregation approaches such as NetVLAD [9, 19, 2], which computes the residuals w.r.t. to feature clusters, generalized-mean pooling (GeM) [13] or global max pooling (MAC) [20] which operate directly on the feature space. Inspired by these works, we propose an aggregation approach that combines, in a single descriptor, a first-order and a second-order aggregation approach. Specifically, we combine the global average pooling aggregator with a pairwise feature interaction aggregator into a single descriptor. We compare our hypothesis with the SOTA models LOGG3D-Net [10], PointNetVLAD [9], and OverlapTransformer [8], and show through results on real-world datasets that the combination of these two statistical-based aggregators yields higher performance than each aggregator individually and also the SOTA models.
@@ II-B Datasets
Less studied is 3D LiDAR place recognition in field or agricultural environments. One possible reason for this is the lack of available datasets which, recently, has been partially compensated by datasets such as Wild-places [25], RELLIS-3D [26], ORFD [27], TreeScope [6], or VineLiDAR [28], showing the increasing interest in such environments. Thus, given the lack of real-world data, especially in horticulture, we propose a novel dataset HORTO-3DLM that includes 3D LiDAR data of semi-structured horticulture environments, particularly from fruit trees namely apples and strawberries. The goal of HORTO-3DLM is to contribute with real-world quality data suitable to place recognition/localization in the domains related to field and agricultural robotics.
@@ III-B2 Feature Aggregation (T2)
where $f_{j}\in\mathbb{R}^{c}$ represents the $j$-th feature point, and $\bar{F}\in\mathbb{R}^{c}$ represents the global average representation. As for the aggregator based on the pairwise feature interactions, it is computed using the outer product of $F$ as follows:
$G=\frac{1}{n}F\,F^{T},$ (2)
where $G\in\mathbb{R}^{c\times c}$ represents pairwise feature interactions of the local features, which is flattened to form a vector representation $G^{\prime}\in\mathbb{R}^{c^{2}}$. Finally, to create the final descriptor $D$, both representations are concatenated into a single vector $D^{*}=[G^{\prime},\bar{F}]\in\mathbb{R}^{c+c^{2}}$, which is fed to a full-connected layer $\text{FC}:\mathbb{R}^{c+c^{2}}\rightarrow\mathbb{R}^{d}$, and normalized using the $L_{2}$-norm:
$D=L_{2}\text{-norm}(\text{FC}(D^{*})).$ (3)
@@ III-C1 Training Data (B1)
Place recognition methods traditionally resort to contrastive learning techniques, which learn to distinguish between similar (positive pairs) and dissimilar (negative pairs) data samples. In place recognition for urban environments, positive pairs are defined only by a fixed radius, which is a sufficient condition given the scale of the trajectories and revisits in these environments. However, to perform place recognition in crops with narrow rows, the radius is a necessary but not sufficient condition. The search conditions must guarantee that the search occurs along the row and does not transgress into neighboring rows. Splitting the crops into unique segments and confining the search within the segment ensures the necessary conditions.
@@ III-C2 Segment-level consistency model (B2)
The SLC model introduces an additional signal during training, which is more context-related than the one generated by the contrastive loss. The goal is to resolve ambiguities created by scans that are highly overlapped but belong to neighboring segments. The SLC model is framed as a classification problem, aiming to predict the segment class using the descriptors as inputs. The loss between the predictions and the corresponding ground-truth segment class $s\in[1,2,...,L]$, where $L$ is the number of segments in a sequence (as shown in Fig. 3), is computed based on the NLL Loss. Hence, given a training tuple of descriptors $\tau^{D}$, and the corresponding tuple of ground-truth segment classes $\tau^{S}=(s^{A},s^{P},\{s^{N}_{j}\})$, all descriptors $D_{i}\in\tau^{D}$ are fed to an MLP to predict the corresponding segment class, such that:
@@
where $\mathbf{p}_{i}=[p_{1},p_{2},...,p_{L}]$ with $p_{k}\in[0,1]$ representing the segment prediction scores which are used, together with the respective ground-truth segment label $s_{i}\in\tau^{S}$, to compute the loss as follows:
@@ IV EXPERIMENTAL EVALUATION
To support the experiments under realistic and challenging cross-horticultural domains, we introduce a new dataset comprising four sequences from distinct horticultural crops, including detailed annotations for ground truth evaluation. This section outlines the evaluation protocol, implementation details, and training procedures, including preprocessing steps and the architecture of the proposed approach. Additionally, quantitative metrics and qualitative retrieval results are reported, showcasing the effectiveness of the approach. Furthermore, we conduct a runtime analysis to assess the computational efficiency of the PointNetPGAP.
@@ IV-A The HORTO-3DLM Dataset
TABLE I: Summary of the four sequences. The various sequence names represent the first letter of plantation type, month and year of the recording. For example, an orchard sequence recorded in November 2022 is named ON22. The columns have the following meaning: name of the sequence (Sequence), month of recording (M), year of recording (Y), number of rows each sequence comprises (N.Rs), number of segments each sequence comprises (N.Segs), traveled distance (D) and plantation type.
Sequence M Y Total Scans N.Rs N.Segs D [m] Plantation Type
@@
TABLE II: Training and testing statistics used in the experiments.
Cross-validation
Sequence Test Anchors Training Anchors Seq. Training Anchors
OJ22 1797 1605 100
OJ23 4801 1193 512
ON22 4993 1210 495
SJ23 3509 1107 598
IV-B Methodology for Place Recognition
@@
Following the protocol defined in Section III-C1, $r_{th}$ was set to 2 m for the training data, which means that a positive pair exists whenever an anchor-positive pair is within a range of 2 m. Additionally, from all anchor-positive pairs, only the anchors that were at least 0.5 m apart were selected for training. For evaluation purposes, $r_{th}$ was set to 10 m, and all anchors were used.
IV-C Evaluation Protocol
The evaluation of the proposed approach is based on a k-fold cross-validation protocol, with $k=4$, which consists of training on three sequences and testing on the sequence that was left out. This is repeated four times to obtain results on all sequences.
The dataset split for the cross-validation protocol is outlined in Table II. The entries have the following meaning: the first column (Sequence) identifies the sequence on which a given model was tested, meaning the model was trained on the remaining sequences; the second column (Test Anchors) refers to the number of anchors/loops with which the model was tested; the third column (Training Anchors) refers to the number of anchors/loops the model was trained with, corresponding to the sum of the training anchors of the remaining sequences; the fourth column (Seq. Training Anchors) refers to the number of training anchors each sequence contains individually.
As for the performance metrics, the results obtained from the proposed experiments are reported using Recall for the top-k retrieved candidates (i.e., Recall@k).
@@ IV-D Training and Implementation Details
All models were trained with 200 epochs on an NVIDIA GeForce RTX 3090 GPU, using the closest positive and 20 negatives ($m=20$) for each anchor. The margin value of the contrastive loss was set to 0.5 ($m=0.5$), and the model parameters were optimized using the AdamW optimizer with a learning rate ($L_{r}$) of 0.0001 and a weight decay ($W_{d}$) of 0.0005. Moreover, all proposed experiments were conducted on Python 3.8, and PyTorch with CUDA 11.6.
@@ IV-E Retrieval Evaluation and Discussion
This section presents and discusses the empirical results. Firstly, we conduct an ablation study on the proposed PointNetPGAP model, delineating the contribution of each aggregator to the performance. Secondly, we benchmark PointNetPGAP, both trained with and without the SLC model, against SOTA models: PointNetVLAD [9], LOGG3D-Net [10], and OverlapTransformer [8]. Each of these models adopts distinct strategies for handling scans: PointNetVLAD processes scans directly, OverlapTransformer processes scans as range images, and LOGG3D-Net relies on a voxel-based representation and network. The main results are presented in Table III and Fig. 4, while results at the segment-level are presented in Fig. 3. Furthermore, a qualitative assessment is outlined in Fig. 6. Finally, the efficiency of the models is evaluated through runtime analysis. Throughout this section, PointNetPGAP will be abbreviated as PNPGAP, and PNPGAP-SLC when trained without and with the SLC model, respectively.

### v2

@@ Abstract
3D LiDAR-based place recognition remains largely underexplored in horticultural environments, which present unique challenges due to their semi-permeable nature to laser beams. This characteristic often results in highly similar LiDAR scans from adjacent rows, leading to descriptor ambiguity and, consequently, compromised retrieval performance. In this work, we address the challenges of 3D LiDAR place recognition in horticultural environments, particularly focusing on inter-row ambiguity by introducing three key contributions: (i) a novel model, PointNetPGAP, which combines the outputs of two statistically-inspired aggregators into a single descriptor; (ii) a Segment-Level Consistency (SLC) model, used exclusively during training to enhance descriptor robustness; and (iii) the HORTO-3DLM dataset, comprising LiDAR sequences from orchards and strawberry fields. Experimental evaluations conducted on the HORTO-3DLM and KITTI Odometry datasets demonstrate that PointNetPGAP outperforms state-of-the-art models, including OverlapTransformer and PointNetVLAD, particularly when the SLC model is applied. These results underscore the model’s superiority, especially in horticultural environments, by significantly improving retrieval performance in segments with higher ambiguity. The dataset and the code will be made publicly available at https://github.com/Cybonic/PointNetPGAP-SLC.git
@@ I INTRODUCTION
In recent years, significant advancements have been made in 3D LiDAR-based place recognition, driven primarily by the autonomous driving community with the aim of achieving long-term localization [1, 2]. While initial efforts have been made to adapt these methods for agricultural environments [3, 4], several challenges unique to these settings remain largely unresolved. Unlike urban environments, where LiDAR scans capture well-defined geometries (e.g., corners, planes), horticultural environments produce scans with sparse and poorly defined features. The semi-permeable nature of these environments allows laser beams to traverse multiple row layers, resulting in scans that cover several rows. This phenomenon leads to sparse scans that overlap with scans from adjacent rows, thereby causing both intra- and inter-row descriptor ambiguity.
This work addresses the challenges of 3D LiDAR place recognition in horticultural environments, particularly focusing on inter-row ambiguity, by introducing three key contributions: (i) PointNetPGAP: A lightweight model that combines PointNet for local feature extraction and integrates the outputs of two statistically-inspired aggregation approaches into a single descriptor. (ii) Segment-Level Consistency (SLC) Model: A training model designed to reduce inter-row ambiguity by differentiating scans from different segments. (iii) HORTO-3DLM Dataset: A novel dataset collected using a 3D LiDAR sensor, comprising data from three orchards and a strawberry plantation.
We benchmark PointNetPGAP at both the sequence and segment levels against state-of-the-art (SOTA) models, specifically OverlapTransformer [5], PointNetVLAD [6], and LOGG3D-Net [7], using the HORTO-3DLM dataset within a cross-validation protocol. Additionally, we evaluate the cross-domain generalization capabilities using the KITTI Odometry dataset [8].
The results highlight that the aggregation approach of PointNetPGAP is better suited to horticultural environments compared to existing SOTA models. Although its impact is reduced in urban settings, it nonetheless exhibits superior rotation invariance in urban environments compared to SOTA models. Moreover, when models are trained with the SLC approach, results at both the sequence and segment levels indicate improved , particularly by improving retrieval accuracy in segments with higher ambiguity.
@@
Figure 2: Top: The retrieval framework uses PointNetPGAP to generate global descriptors through the following stages: T1) PointNet processes an input scan $P$ to extract local features $F$; T2) These features are aggregated into a global descriptor $D$ by combining global average representation with pairwise feature interactions; T3) The global descriptor is then used to query the database and retrieve the top-$k$ most similar places. Bottom: The training scheme involves: B1) Inputting a training tuple with an anchor-positive pair and $m$ negatives to generate descriptors; B2) Feeding these descriptors into an MLP in the SLC model to predict segment classes, the SLC loss computed via Negative Log-likelihood using the predictions and the segment labels; B3) Calculating the LazyTriplet loss based on anchor-positive and anchor-negative distances; B4) Combining both losses with a weight $\alpha$.
@@ II RELATED WORK
Place recognition has been the focus of extensive research in recent years, particularly within the autonomous driving community [1], where Deep Learning-based approaches have garnered the most attention. One potential reason for the limited research in other domains may be the scarcity of publicly available datasets.
@@
Regarding aggregation strategies, most 3D LiDAR place recognition frameworks rely on first-order aggregation approaches such as NetVLAD [6, 18, 14], which computes the residuals w.r.t. feature clusters, generalized-mean pooling (GeM) [11] or global max pooling (MAC) [19] which operate directly on the feature space. Inspired by these works, we propose an aggregation approach that combines, in a single descriptor, a first-order and a second-order aggregation approach. Specifically, we combine the global average pooling aggregator with a pairwise feature interaction aggregator into a single descriptor. We compare our hypothesis with the SOTA models LOGG3D-Net [7], PointNetVLAD [6], and OverlapTransformer [5], and show through results on real-world datasets that the combination of these two statistical-based aggregators yields higher performance than each aggregator individually and also the SOTA models.
@@ II-B Datasets
Less studied is 3D LiDAR place recognition in field or agricultural environments. One possible reason for this is the lack of available datasets which, recently, has been partially compensated by datasets such as Wild-places [23], RELLIS-3D [24], ORFD [25], TreeScope [4], or VineLiDAR [26], showing the increasing interest in such environments.
@@ III-B2 Feature Aggregation (T2)
where $f_{j}\in\mathbb{R}^{c}$ represents the $j$-th feature vector, and $\bar{F}\in\mathbb{R}^{c}$ represents the global average representation. As for the aggregator based on the pairwise feature interactions, it is computed as follows:
$G=\frac{1}{n}F^{T}\,F,$ (2)
where $G\in\mathbb{R}^{c\times c}$ represents pairwise feature interactions of the local features, which is flattened into a vector representation $G^{\prime}\in\mathbb{R}^{c^{2}}$. Finally, to create the final descriptor $D$, both representations are concatenated into a single vector $D^{*}=[G^{\prime},\bar{F}]\in\mathbb{R}^{c+c^{2}}$, which is fed to a full-connected layer $\text{FC}:\mathbb{R}^{c+c^{2}}\rightarrow\mathbb{R}^{d}$, and normalized using the $L_{2}$-norm: $D=L_{2}\text{-norm}(\text{FC}(D^{*}))$.
@@ III-C1 Training Data (B1)
Place recognition methods traditionally resort to contrastive learning techniques, which learn to distinguish between similar (positive pairs) and dissimilar (negative pairs) data samples. In place recognition for urban environments, positive pairs are defined only by a fixed radius, which is a sufficient condition given the scale of the trajectories and revisits in these environments. However, to perform place recognition in crops with narrow rows, the radius is a necessary but not sufficient condition. The search conditions must guarantee that the search occurs along the row and does not transgress into neighboring rows. Dividing the crops into distinct segments and confining the search within each segment ensures the necessary conditions.
@@ III-C2 Segment-level consistency model (B2)
The SLC model introduces an additional signal during training, which is more context-related than the one generated by the contrastive loss. The goal is to resolve ambiguities created by scans that are highly overlapped but belong to neighboring segments. The SLC model is framed as a classification problem, aiming to predict the segment class using the descriptors as inputs. The loss between the predictions and the corresponding ground-truth segment class $s\in\{1,2,...,L\}$, where $L$ is the number of segments in a sequence (as shown in Fig. 3), is computed based on the NLL Loss. Hence, given a training tuple of descriptors $\tau^{D}$, and the corresponding tuple of ground-truth segment classes $\tau^{S}=(s^{A},s^{P},\{s^{N}_{j}\})$, all descriptors $D_{i}\in\tau^{D}$ are fed to an MLP to predict the corresponding segment class, such that:
@@
where $\mathbf{p}_{i}=[p_{1},p_{2},...,p_{L}]^{T}$ with $\sum\limits_{p_{k}\in\mathbf{p}_{i}}p_{k}=1$ representing the segment prediction likelihood. The predictions are used, together with the respective ground-truth segment label $s_{i}\in\tau^{S}$, to compute the loss as follows:
@@ IV EXPERIMENTAL EVALUATION
This section outlines the evaluation protocol, implementation details, and training procedures of the proposed approach. Both quantitative and qualitative retrieval results are presented to support the proposed contributions. Additionally, a runtime analysis is performed to assess the computational efficiency of PointNetPGAP.
@@ IV-A The HORTO-3DLM Dataset
TABLE I: Summary of the four sequences. The various sequence names represent the first letter of plantation type, month and year of the recording. For example, an orchard sequence recorded in November 2022 is named ON22. The columns have the following meaning: name of the sequence (Seq), month of recording (M), year of recording (Y), number of rows each sequence comprises (N.Rs), number of segments each sequence comprises (N.Segs), traveled distance (D) and plantation type.
Seq. M Y NºScans N.Rs N.Segs D [m] Plantation Type
@@
TABLE II: Cross-Validation training and evaluation data. Each column represents the training and test data of the evaluated sequence (row 1)
OJ22 OJ23 ON22 SJ23
Nº Training Anchors ($r_{th}\leq 2m$) 1605 4801 1210 1107
Nº Test Anchors ($r_{th}\leq 10\text{m}$) 1797 1193 4993 3509
TABLE III: KITTI Odometry dataset and cross-domain training and evaluation data
  KITTI Ondemtry dataset 00 02 05 06 08 Nºof Frames 4051 4661 2761 1101 4071 Nº of Anchors ($r_{th}\leq 10\text{m}$) 809 344 454 273 353 Cross Domain Training / Evaluation data Sequences Nº Anchors Training OJ22, OJ23, SJ23 ON22 1705 Validation 00 809 Test 02, 05, 06, 08 1424
IV-B HORTO-3DLM Dataset for Place Recognition
@@
Following the protocol defined in Section III-C1, $r_{th}$ was set to 2 m for the training data, which means that a positive pair exists whenever an anchor-positive pair is within a range of 2 m. Additionally, from all training anchor-positive pairs, only the anchors that are at least 0.5 m apart were selected for training.
IV-C Evaluation Protocols
The proposed approach is evaluated based on two strategies: a 4-fold cross-validation protocol; and a cross-domain approach. Performance metrics are reported using Recall for the top-k retrieved candidates (i.e., Recall@k).
IV-C1 Cross-Validation
This protocol evaluates generalization capabilities within the same domain. The HORTO-3DLM dataset is split into four sequences, with models trained on three sequences and evaluated on the remaining one. This process is repeated four times to cover all sequences. The dataset split is outlined in Table II. A retrieved candidate is considered a true positive if it is within an $10$ m radius of the query and belongs to the same segment; otherwise, it is considered a false positive.
IV-C2 Cross-Domain
This protocol assesses generalization capabilities across domains, specifically from horticultural to urban environments. The models are trained on the HORTO-3DLM dataset and evaluated on the KITTI Odometry dataset [8]. The data split regarding this evaluation strategy is presented in Table III. A retrieved candidate is considered a true positive if it is within an $10$ m radius of the query (no segments are considered); otherwise, it is considered a false positive.
@@ IV-D Training and Implementation Details
All models were trained with 200 epochs, using early stopping to obtain the best performance on the validation set. All models were trained on an NVIDIA GeForce RTX 3090 GPU, using the closest positive and 20 negatives ($m=20$) for each anchor. The margin value of the contrastive loss was set to 0.5 ($m=0.5$), and the model parameters were optimized using the AdamW optimizer with a learning rate ($L_{r}$) of 0.0001 and a weight decay ($W_{d}$) of 0.0005. Moreover, all proposed experiments were conducted on Python 3.8, and PyTorch with CUDA 11.6.
TABLE IV: Recall@1 performance of the cross-validation protocol (same domain). Reporting the results of the models trained: (without SLC)/(with SLC). Bolded scores indicate the highest value between models trained without and with SLC, while underlined scores represent the highest score in each column (i.e., the best-performing model for each sequence).
@@ IV-E Retrieval Evaluation and Discussion
This section presents and discusses the empirical results of the work. First, an ablation study is conducted on the proposed PointNetPGAP model to showcase the contribution of each aggregator to the overall performance. Subsequently, PointNetPGAP is benchmarked within the same domain, both with and without the SLC model, against SOTA models: PointNetVLAD [6], LOGG3D-Net [7], and OverlapTransformer [5]. These models employ different strategies for processing LiDAR scans: PointNetVLAD processes scans directly, OverlapTransformer processes scans as range images, and LOGG3D-Net utilizes a voxel-based representation. Finally, the generalization capabilities of the proposed approach across domains are evaluated.
Throughout this section, PointNetPGAP is abbreviated as PNPGAP, and PNPGAP-SLC when trained with the SLC model.
TABLE V: Ablation study of PNPGAP, where GAP represents the global average pooling aggregator and PFI represents the aggregator based on pairwise feature interactions. The bolded scores represent the highest values in each column.